OpenStudy (anonymous):

Solve the following system using the substitution method. 10x + y = -1 5x + 2y = 16

OpenStudy (anonymous):

okay take y seperate from 1st equation and put it in 2nd y=-1-10x

OpenStudy (anonymous):

putting it in 2nd equtaion 5x+2(-1-10x)=16

OpenStudy (anonymous):

5x-2-20x=16 -15x=18 x=-18/15 x=-6/5

OpenStudy (anonymous):

now put x in y=-1-10x y=-1-10(-6/5) y=-1+60/5 y=-1+12 y=11, x=-6/5

OpenStudy (anonymous):

ok, thanks again khurram!

OpenStudy (anonymous):

welcome